slope intercept form of 3x-2y=-19
Answer by nerdybill(7384)

slope intercept form of 3x-2y=-19
.
slope-intercept form is
y = mx + b
.
simply move terms around to the above form:
3x-2y=-19
subtracting 3x from both sides:
-2y=-3x-19
multiplying both sides by -1:
2y=3x+19
divide both sides by 2:
y = (3/2)x + 19/2 (this is what they're looking for)